Kahn Inc. has a target capital structure of 70% common equity and 30% debt to fund its $11 billion in operating assets. Furthermore, Kahn Inc. has a WACC of 14%, a before-tax cost of debt of 9%, and a tax rate of 25%. The company's retained earnings are adequate to provide the common equity portion of its capital budget. Its expected dividend next year (D1) is $3, and the current stock price is $35.


 a. What is the company's expected growth rate? Do not round intermediate calculations. Round your answer to two decimal places. 

 b. If the firm's net income is expected to be $1.5 billion, what portion of its net income is the firm expected to pay out as dividends? Do not round intermediate calculations. Round your answer to two decimal places. (Hint: Refer to Equation below.) 


Growth rate = (1 - Payout ratio)ROE
Answer #1

0.14 = 0.70(0.09)(1 - 0.25) + 0.30(Ce)

Ce = 30.92%

a.

Using Constant Growth Model,

g = 0.3092 - 3/35

g = 22.35%

b.

ROE = 1.5/(0.30*11)

ROE = 45.45%

0.2235 = (1 - b)(0.4545)

b = 50.83%

Dividend Payout Ratio = 50.83%